Need help solving

Dobbs Company issues 6%, two-year bonds, on December 31, 2017, with a par value of $101,000 and semiannual interest payments. Unamortized Discount Carrying Value (O) 12/31/2017 (1) 6/30/2018 (2) 12/31/2018 (3) 6/30/2019 (4) 12/31/2019 $6,020 4, 515 3,010 1,505 94,980 96,485 97,990 99,495 101,000 Use the above straight-line bond amortization table and prepare journal entries for the following. Required: (a) The issuance of bonds on December 31, 2017 (b) The first through fourth interest payments on each June 30 and December 31. (c) Record the maturity of the bonds on December 31, 2019.
media%2Ff2a%2Ff2aceaff-c71e-4fb8-a9cb-0a
media%2F5ca%2F5cabf4a2-395f-4b22-a7b1-81
media%2F475%2F475df306-75ab-4f1d-a45b-9c
Answer #1

Solution

Dobbs Company

Journal Entries:

Dobbs Company

Date

Account Titles and Explanation

Debit

Credit

31-Dec-17

Cash

$94,980

Discount on Bonds Payable

$6,020

Bonds Payable

$101,000

(To record issuance of bonds at discount)

30-Jun-18

Bond Interest Expense

$4,535

Discount on Bonds Payable

$1,505

Cash

$3,030

(To record bond interest expense and amortization of discount on bonds payable)

31-Dec-18

Bond Interest Expense

$4,535

Discount on Bonds Payable

$1,505

Cash

$3,030

(To record bond interest expense and amortization of discount on bonds payable)

30-Jun-19

Bond Interest Expense

$4,535

Discount on Bonds Payable

$1,505

Cash

$3,030

(To record bond interest expense and amortization of discount on bonds payable)

31-Dec-19

Bond Interest Expense

$4,535

Discount on Bonds Payable

$1,505

Cash

$3,030

(To record bond interest expense and amortization of discount on bonds payable)

31-Dec-19

Bonds Payable

$101,000

Cash

$101,000

(To record bonds payable redeemed on maturity)

Computations:

  1. Semiannual interest expense at 6% -

$101,000 x 6% x 6/12 = $3,030

  1. Amortization of discount on bonds payable –

Total Discount on bonds payable = 101,000 – 94,980 = $6,020

Amortization of discount on bonds payable on semiannual interest dates = 6,020/4 = $1,505

The period to maturity is two years. With two semiannual periods the company makes 4 (2 x 2) interest payments and hence amortization of discount on bonds is computed for those days.

  1. Interest expense = cash paid + amortization of discount on bonds payable

= $3,030 + $1,505 = $4535